A bacteria culture starts with 120 bacteria and after three hours the population consists of 200 bacteria. assuming that the culture grows exponentially, find the poplation after 5 hours.
Answer by stanbon(75887)

A bacteria culture starts with 120 bacteria and after three hours the population consists of 200 bacteria. assuming that the culture grows exponentially, find the poplation after 5 hours.
---------------------------------------------------
A(t) = ab^t
---------------------
200 = 120b^3
b^3 = 200/120 = 5/3
b = 1.047197551....
-----------
Equation:
A(5) = 120(1.047197551)^5
A(5) = 151.12
Rounding down, A(5) = 151
=============================
Cheers,
Stan H.